OpenStudy (summersnow8):

12-15: on the average 20% of seniors develop influenza during a given year. 15 students are selected at random from a nursing home & followed up for a year. P= .2 1-P= .8 12. What is the probability exactly 3 out of 15 seniors will develop influenza? 13. What is the probability that at most 3 of them will develop influenza? 14. What is the probability that at least 3 of them will develop influenza? 15. The standard deviation of the number of seniors will get influenza is: please show work so I can understand how to solve :) thanks

OpenStudy (kropot72):

@Summersnow8 Have you learned the binomial distribution yet?

OpenStudy (summersnow8):

yes, but I am confused about how to solve this, my professor is terrible at explaining things

OpenStudy (summersnow8):

I know how to use the binompdf or binomcdf on my calculator, but can't seem to figure this problem out.

OpenStudy (dumbcow):

std dev = sqrt(n*p*(1-p))

OpenStudy (kropot72):

12 can be solved as follows: \[P(3\ out\ of\ 15)=15C3\times(0.2)^{3} \times(0.8)^{12}=you\ can\ calculate\]

OpenStudy (summersnow8):

15 nPr 3 x....etc?

OpenStudy (summersnow8):

what does the C stand for? nCr?

OpenStudy (summersnow8):

I got .25 for 12

OpenStudy (kropot72):

15C3 means the number of combinations of 15 different items taken 3 at a time. Your answer for 12 is correct.

OpenStudy (dumbcow):

For 13 find probabilities for x=0,1,2,3 add them up to get P(X<= 3) for 14 sum x=0,1,2 subtract this sum from 1 this will give P(X>=3)

OpenStudy (summersnow8):

ok how do you do the rest of the problems @kropot72

OpenStudy (summersnow8):

@dumbcow that doesn't help me understand how to solve the problem

OpenStudy (dumbcow):

ok let me explain "x" is the number out of 15 that get sick, from 0 to 15 the sum of all probabilities is always 1 you want the probability that "at most" 3 get sick, so you consider the cases where 0,1,2,3 people out of 15 get sick using the formula @kropot72 showed above to get individual probabilities for each case add them up and that gives you P(x<= 3) or probability that at most 3 get sick

OpenStudy (kropot72):

Going over the method of calculation for the binomial distribution: If we let p stand for the probability of success, and q = 1 - p the probability of failure, then the probability of r successes when n independent trials are carried out is given by: \[\left(\begin{matrix}n \\ r\end{matrix}\right) \times p ^{r} \times (1-p)^{n-r}\] where the binomial coefficient \[\left(\begin{matrix}n \\ r\end{matrix}\right)=\frac{n!}{r!(n-r)!}=nCr\]

OpenStudy (kropot72):

So to find the probability of zero getting sick, we substitute the following values: n = 15 r = 0 p = 0.2 1 - p = 0.8 Can you now find P(0 out of 15) ?

OpenStudy (summersnow8):

i got .035

OpenStudy (kropot72):

Good work! Your calculation is correct. Now calculate for r = 1, r = 2 and r = 3 as @dumbcow stated.

OpenStudy (summersnow8):

what number would that be for?

OpenStudy (kropot72):

The sample size n = 15 for each calculation.

OpenStudy (summersnow8):

I don't understand how that helps me solve 13-15

OpenStudy (summersnow8):

is there a faster way instead of calculating each? because how does it differ from AT LEAST 3 and AT MOST 3?

OpenStudy (summersnow8):

@kropot72?

OpenStudy (summersnow8):

because I added them all up and I got 5.28....

OpenStudy (kropot72):

@Summersnow8 Sorry, I will be tied up on another matter for about an hour. Can help again later. :)

OpenStudy (summersnow8):

@kropot72 oh okay

OpenStudy (kropot72):

We already found some of the values of probability: p(0 of 15 infected) = 0.0352 p(1 of 15 infected) = ? p(2 of 15 infected) = ? p(3 of 15 infected) = 0.25 What values did you get for p(1 of 15 infected) and p(2 of 15 infected) ?

OpenStudy (ybarrap):

12. What is the probability exactly 3 out of 15 seniors will develop influenza? $$ \large{ \binom{15}{3}.2^3.8^{12} }\\ =0.25013889531904 $$ http://www.wolframalpha.com/input/?i=%2815choose3%29*.2^3*.8^12 13. What is the probability that at most 3 of them will develop influenza? $$ \large{ P(X \le 3) = P(X=0)+P(X=1)+P(X=2)+P(X=3)\\ =\binom{15}{0}.2^0.8^{15}+\binom{15}{1}.2^1.8^{14}+\binom{15}{2}.2^2.8^{13}+\binom{15}{3}.2^3.8^{12} }\\ =0.648162104573952 $$ http://www.wolframalpha.com/input/?i=%2815choose0%29*.2^0*.8^15%2B%2815choose1%29*.2^1*.8^14%2B%2815choose2%29*.2^2*.8^13%2B%2815choose3%29*.2^3*.8^12 14. What is the probability that at least 3 of them will develop influenza? $$ P(X\ge3)=1-P(X<3)=1-P(X\le2)\\ =1-\left [\binom{15}{0}.2^0.8^{15}+\binom{15}{1}.2^1.8^{14}+\binom{15}{2}.2^2.8^{13}\right]\\ =0.6019767907450880 $$ http://www.wolframalpha.com/input/?i=1-%28%2815choose0%29*.2^0*.8^15%2B%2815choose1%29*.2^1*.8^14%2B%2815choose2%29*.2^2*.8^13%29 Does this make sense?

OpenStudy (ybarrap):

15. The standard deviation, \(\sigma\), of the number of seniors will get influenza is: $$ \sigma^2=npq=15\times.2\times.8=2.4\\ \sigma=\sqrt{2.4}=1.54919... $$ http://en.wikipedia.org/wiki/Binomial_distribution#Mean_and_variance Hope this helped.

OpenStudy (summersnow8):

@ybarrap perfect, thank you